# FlowMAC: Conditional Flow Matching for Audio Coding at Low Bit Rates

### v1

@@ I Introduction
In the modern digital world, audio codecs are used on a day-to-day basis, so every technological advancement can have a large impact. In recent years, deep neural networks (DNNs) revolutionized the field of audio compression. Early approaches [1, 2, 3] control the compression at via entropy-based losses and ensure good quality via reconstruction losses. With the advent of deep generative models the quality of neural codecs at bit rates lower than 12 kbps greatly improved.
@@ II Flow Matching fundamentals
where $\mathbf{x}_{1}\sim q(\mathbf{x}_{1})$ sampled from the train set, $\mu_{t}(x)=t\mathbf{x}_{1}$, and $\sigma_{t}=1-(1-\sigma_{\text{min}})t$ with $\sigma_{\text{min}}\ll 1$. This defines a Gaussian path where $p_{0}$ is the standard Gaussian and $p_{1}$ is a Gaussian centered at $\mathbf{x}_{1}$ with small variance. Theorem 3 in [19] shows that this probability path is generated by the Optimal Transport Conditional Vector Field
@@ III-D FlowMAC inference
Early experimentation showed that the quality of the mel coder subsystem quickly saturates. To test this, we introduce FlowMAC-CQ, a separately trained model at 6 kbps. For this we use the same hyperparameters and NFE as for FlowMAC. Finally, we test the quality-complexity trade-off via using a single step for the Euler method and no CFG, hence, obtaining FlowMAC-LC and using 1 NFE.
@@ V Results and Discussion
Figure 2 illustrates the results of the P.808 DCR listening test with 46 listeners. The results from the naive listeners confirm that both FlowMAC and FlowMAC-LC are the best models at 3 kbps, being on average on par with EnCodec and MBD at 6 kbps. FlowMAC at 1.5 kbps shows a significant quality drop, while no significant quality improvement is achieved by the 6 kbps version FlowMAC-CQ. As expected, the copy-synthesis with BigVGAN offers the highest achievable quality for our system. FlowMAC-LC’s average rating are lower than the high-complexity version. Still, the test confirms that it is a competitive baseline. For naive listeners the higher frequency resolution of DAC 44.1 kHz at 8 kbps offers does not offer an advantage over the 24 kHz model.
Overall we notice that the all DNN conditions achieve comparable quality with the legacy USAC condition at similar bit rates, the only exception being FlowMAC at 3 kbps.
The results of the MUSHRA test with 14 listeners are illustrated in Figure 3. While USAC 8 kbps has a quality advantage on average over the other codecs here, this test demonstrates that FlowMAC at 3 kbps performs similar to DAC 6 kbps and both conditions outperform MBD 6 kbps. We notice that the performance of the DNN-based codecs highly varies for different items in the test set. In particular, FlowMAC performs poorly on the out-of-distribution test items, while its performance is on average comparable with DAC for speech and music. The copy-synthesis from BigVGAN performs best average and offers an measurement of the highest quality achievable with FlowMAC. We notice that these results more clearly highlight fine difference between the codecs, but are overall in accordance with our P.808 test results.

### v2

@@ I Introduction
In the modern digital world, audio codecs are used on a day-to-day basis, so every technological advancement can have a large impact. In recent years, deep neural networks (DNNs) revolutionized the field of audio compression. Early approaches [1, 2, 3] control the compression via entropy-based losses and ensure good quality via reconstruction losses. With the advent of deep generative models the quality of neural codecs at bit rates lower than 12 kbps greatly improved.
@@ II Flow Matching fundamentals
where $\mathbf{x}_{1}\sim q(\mathbf{x}_{1})$ sampled from the train set, $\mu_{t}(\mathbf{x}_{1})=t\mathbf{x}_{1}$, and $\sigma_{t}(\mathbf{x}_{1})=1-(1-\sigma_{\text{min}})t$ with $\sigma_{\text{min}}\ll 1$. This defines a Gaussian path where $p_{0}$ is the standard Gaussian and $p_{1}$ is a Gaussian centered at $\mathbf{x}_{1}$ with small variance. Theorem 3 in [19] shows that this probability path is generated by the Optimal Transport Conditional Vector Field
@@ III-D FlowMAC inference
Early experimentation showed that the quality of the mel coder subsystem quickly saturates. To test this, we introduce FlowMAC-CQ, a separately trained model at 6 kbps. For this we use the same hyperparameters and NFE as for FlowMAC. Finally, we test the quality-complexity trade-off via using a single step for the Euler method and no CFG, hence, obtaining FlowMAC-LC (Low Complexity) and using 1 NFE.
@@ V Results and Discussion
Figure 2 illustrates the results of the P.808 DCR listening test with 46 listeners. The results from the naive listeners confirm that both FlowMAC and FlowMAC-LC are the best models at 3 kbps, being on average on par with EnCodec and MBD at 6 kbps. FlowMAC at 1.5 kbps shows a significant quality drop, while no significant quality improvement is achieved by the 6 kbps version FlowMAC-CQ. As expected, the copy-synthesis with BigVGAN offers the highest achievable quality for our system. FlowMAC-LC’s average rating are lower than the high-complexity version. Still, the test confirms that it is a competitive baseline. For naive listeners the higher frequency resolution of DAC 44.1 kHz at 8 kbps does not offer an advantage over the 24 kHz model.
Overall we notice that all DNN conditions achieve comparable quality with the legacy USAC condition at similar bit rates, the only exception being FlowMAC at 3 kbps.
The results of the MUSHRA test with 14 listeners are illustrated in Figure 3. While USAC 8 kbps has a quality advantage on average over the other codecs here, this test demonstrates that FlowMAC at 3 kbps performs similar to DAC 6 kbps and both conditions outperform MBD 6 kbps. We notice that the performance of the DNN-based codecs highly varies for different items in the test set. In particular, FlowMAC performs poorly on the out-of-distribution test items, while its performance is on average comparable with DAC for speech and music. The copy-synthesis from BigVGAN performs best average and offers a measurement of the highest quality achievable with FlowMAC. We notice that these results more clearly highlight fine difference between the codecs, but are overall in accordance with our P.808 test results. [^1] Check our demo samples at: https://fhgspco.github.io/flowmac